OpenStudy (barrelracing):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form. a) y=3x + 2 b) y= 3x + 8 c) y = -3x - 2 d) y = -3x + 8

OpenStudy (barrelracing):

I just need the formula. I cant find it anywhere.

OpenStudy (teddyiswatshecallsme):

Do you know how to calculate the rate of change?

OpenStudy (anonymous):

Use \( y - y_1 = m(x - x_1) \) and then turn it into point slope form

OpenStudy (marihelenh):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] will give you the slope, and then use the formula @Nixy gave you.

OpenStudy (anonymous):

You can get the slope with \( \huge m = \frac{y_2-y_1}{x_2 - x_1} \)

OpenStudy (barrelracing):

y = -3x + 8

OpenStudy (anonymous):

You are correct @Barrelracing